1.   A 1-kg body of velocity 10m/s collides with a 2-kg body which is the rest .After collision, the two stick together. What is the velocity of the two bodies after collision?

 


1
Expert's answer
2022-06-13T08:23:00-0400

Given:

"m_1=1\:\rm kg"

"m_2=2\:\rm kg"

"v_1=10\:\rm m/s"


The law of conservation of momentum says

"m_1v_1=(m_1+m_2)u"

Hence, the final speed of bodies

"u=\frac{m_1v_1}{m_1+m_2}=\frac{1*10}{1+2}=3.3\:\rm m/s"